OpenStudy (anonymous):

Help please! Write the following inequality in slope-intercept form. -8x + 4y ≥ -52 A) y ≥ 2x - 13 B) y ≤ 2x - 13 C) y ≤ 2x + 13 D) y ≥ 2x + 13

OpenStudy (anonymous):

add 8x to both sides, divide both sides by 4. and you are done after folowing these two steps.

OpenStudy (anonymous):

add 8 to the -52?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

-8x+y+8x = -52+8x

OpenStudy (anonymous):

am I right?

OpenStudy (anonymous):

@dinosour1480

OpenStudy (anonymous):

almost

OpenStudy (anonymous):

divide 4y and 52 by 4

OpenStudy (anonymous):

1 & -13

OpenStudy (anonymous):

@dinosour1480

OpenStudy (anonymous):

@LynFran

OpenStudy (lynfran):

\[-8x+4y \ge-52\]\[-8x+8x+4y \ge -52+8x\]\[4y \ge-52+8x\]\[y \ge \frac{ -52 }{ 4 }+\frac{ 8x }{ 4 }\]

OpenStudy (anonymous):

the answer is A? y ≥ 2x - 13

OpenStudy (lynfran):

can also be written as \[y \ge \frac{ 8x }{ 4 }-\frac{ 52 }{ 4 }\]

OpenStudy (lynfran):

yes \[y \ge 2x-13\]

OpenStudy (anonymous):

thank you! :)